# SortWaste: A Densely Annotated Dataset for Object Detection in Industrial Waste Sorting

### v1

@@ 1 Introduction
The exponential growth of the global population and industrial development has increased waste production, making it one of the most pressing environmental challenges of the 21st century. Inefficient waste management systems negatively impact human health, ecosystems, water quality, and the environment at large [b1]. As a result, developing innovative and more effective ways to manage and reduce waste is crucial to building a more sustainable future. Solid waste management is a key challenge, with global generation estimated at 2.01 billion tons annually and projected to reach 3.40 billion tons by 2050 under current trends [b2].
In most European countries, two principal systems are used for waste sorting. The first is a single-stream system, known as Mechanical-Biological Treatment (MBT), which processes Municipal Solid Waste (MSW). The second is a multi-stream system in which materials are selectively collected through source separation, such as plastics. The main challenge for the single-stream system is that valuable recyclables, particularly plastics, are often incorrectly discarded, limiting their recovery through proper sorting. This loss increases landfill use and consumes raw resources, sabotaging sustainability goals. However, as noted in [b3], MBT systems in urban areas can sometimes achieve recovery rates comparable to selective collection. This suggests that with improvements, MBT sorting could be highly impactful. Building on these challenges, deep learning and computer vision offer new opportunities for waste detection and recycling, improving efficiency, reducing reliance on manual labor, and mitigating health risks for workers exposed to hazardous materials. Nevertheless, automatic waste sorting remains challenging because real-world waste streams often contain objects that are deformed, broken, occluded, overlapping, or contaminated [b13].
Among the reasons for the lack of automation in waste sorting is the scarcity of publicly available datasets. Existing public datasets typically do not reflect the complex environments found in real-world waste sorting facilities [b4, b5, b6, b7, b8, b9, b10, b11, b12], except one contribution [b13]. While some companies have developed proprietary datasets for internal use [b26, b27, b28], these datasets are not publicly released, restricting the scientific community’s ability to advance research in this area.
@@ 2.1 Waste Detection Datasets
Litter Detection. The Garbage in Photos (GINI) dataset [b5] contains images of a single class. It was created using the Bing Image Search API and comprises approximately 1,400 images. Similarly, the WADE-AI dataset [b6] includes pictures collected from Google Street View, providing a diverse visual set, though it is also limited to a single generic waste classification. Another relevant dataset is TACO (Trash Annotations in Context) [b7], introduced by Proença et al., designed for waste detection and segmentation. It contains 1,500 mobile-acquired images annotated into 28 waste categories and covers diverse real-world contexts, including outdoor environments, making it valuable to unconstrained applications. OpenLitterMap [b9] is one of the largest public litter datasets, with over 100,000 user-contributed mobile images from around the world, providing diverse scenes and photographic styles that enhance its suitability for real-world litter detection.
@@
Marine Debris. The J-EDI dataset [b14] consists of 5,720 images grouped into three classes, created to train deep learning models that enable autonomous underwater vehicles (AUVs) to detect waste in aquatic environments. In addition, AquaTrash, presented in [b15] and derived from the TACO dataset, contains 369 multi-class annotated images. This dataset was developed to support AquaVision, a deep learning model aimed at detecting and classifying pollutants in the ocean and along the seashore, contributing to water pollution monitoring and mitigation.
@@
General Waste Objects. The Waste Pictures dataset [b8] was constructed using images collected through Google searches. It contains approximately 24,000 images, organized into 34 categories representing a wide variety of waste types. The Taiwan Recycled Waste Dataset (TRWD) [b12] comprises 6,233 images featuring multiple objects and is specifically tailored to the unique waste characteristics found in Taiwan. The dataset is divided into six classes.
@@
Recycling Waste. Thung et al. introduced TrashNet [b4], a widely used benchmark for waste classification comprising 2,400 images in six categories: metal, paper, plastic, glass, organic, and other, captured against a white background under varying lighting. The WaDaBa dataset [b10] focuses exclusively on domestic plastic waste, with 100 plastic objects photographed 40 times each under different conditions, making it particularly relevant for controlled plastic waste identification. Labeled Waste in the Wild (LWW), introduced by Sousa et al. in [b11], is a custom dataset containing 1,000 images of waste in food trays captured under real-world conditions. ZeroWaste-f [b13] is one of the first high-quality waste datasets collected in real-world conditions. Acquired at a Material Recovery Facility (MRF) in the USA from a high-quality paper sorting line, it is split into three parts for different learning paradigms; the supervised ZeroWaste-f subset contains 4,503 annotated images with four classes. While numerous datasets exist for waste detection tasks, most are not specifically designed for industrial applications, except for the ZeroWaste dataset [b13]. Additionally, many datasets consist of images of a single object on a plain white background, which limits the generalizability and effectiveness of models trained on such data when applied to real-world environments. Table 1 summarizes the reviewed datasets, and Figure 2 shows representative examples.
@@
TACO [b7] 1500 28 Segmentation Mask R-CNN Litter detection
WADE-AI [b6] 1396 1
@@
OpenLitterMap [b9] 100000+ 100+ Classification - Litter detection
GINI[b5] 1400 1 Classification - Litter detection
@@
AquaTrash [b15] 369 4 Detection Faster R-CNN, RetinaNet Marine debris
@@
TRWD [b12] 6233 6 Detection YOLOv3 Waste detection
Waste Pictures [b8] 23633 34 Classification - Trash objects
@@
TrashNet [b4] 2400 6 Classification SVM Recycling waste
WaDaBa [b10] 4000 6 Classification - Recycling waste
LWW [b11] 1002 19 Detection Faster R-CNN Recycling waste
ZeroWaste-f [b13] 4503 4
@@ 2.2 Object Detection Models
Faster R-CNN [b17] is a two-stage object detector that extends R-CNN [b16] and Fast R-CNN [b18] by introducing a Region Proposal Network (RPN), which replaces hand-crafted methods with a fully convolutional network using anchor boxes across scales and aspect ratios. The generated proposals are refined by a Fast R-CNN [b18] head for classification and bounding-box regression, and shared convolutional features between the RPN and the detector enable an efficient, end-to-end trainable pipeline that balances accuracy and speed.
Mask R-CNN [b18] extends Faster R-CNN [b17] by adding a parallel branch for instance segmentation: for each Region of Interest (RoI), it predicts a binary mask alongside the class label and bounding box, enabling joint object detection and pixel-level instance segmentation.
Trident Network [b19] addresses the challenge of scale variation in object detection, which often limits the performance of models such as Faster R-CNN [b17] and Mask R-CNN [b18]. The key idea is to use trident blocks: a multi-branch architecture with shared parameters but different dilation rates. Each branch operates with a distinct receptive field, allowing the network to capture objects at different scales and perform scale-aware detection without relying on external image or feature pyramids. This design improves the performance of two-stage detectors while maintaining the inference complexity.
RetinaNet [b20], introduced by Lin et al., is a one-stage detector that reaches accuracy comparable to state-of-the-art two-stage models. Its key contribution is Focal Loss, which mitigates the severe foreground–background imbalance in dense detection. The architecture combines a ResNet–FPN backbone with two subnetworks for anchor classification (using Focal Loss) and bounding box regression. This design substantially closes the accuracy and speed gap between one-stage and two-stage detectors.
YOLO family of detectors [b21] formulates object detection as a single-stage regression problem, enabling real-time performance. Successive versions of YOLO, from YOLOv1 to YOLOv10 [b22, b23], added features such as multi-scale detection, new activation functions, and anchor-free designs. The YOLOv11 [b24] further improves efficiency and accuracy with components like the C3k2 block, SPPF module, and C2PSA mechanism, and extends the framework to instance segmentation, pose estimation, and multi-view recognition.
@@ 3.2 Annotation
The frame rate was downsampled to 5 fps to preserve variability while reducing annotation workload. In total, 18 minutes of video were annotated, yielding 5,396 frames with dirty, deformed, broken, and overlapping objects. Annotations were created over approximately two months using CVAT [b29], with each frame requiring, on average, six minutes, depending on scene complexity and object count. This careful annotation process was essential to ensure the quality of the densely labeled dataset. Although there are various types of plastic materials, some are challenging to distinguish using computer vision. Therefore, the annotation focused on simpler and more easily identifiable plastic classes.
@@ 4.1 Implementation Details
We use Faster R-CNN [b17], TridentNet [b19], and RetinaNet [b20], initialized with weights pretrained on the COCO dataset. For each model, only the learning rate and optimizer are tuned, while the batch size is fixed at eight. Training employs early stopping with a patience of 15 epochs or a maximum of 80,000 iterations. For YOLOv11 [b24], we adopt a similar tuning strategy, focusing on the learning rate and optimizer, with early stopping defined as 15 epochs of patience or a maximum of 300 epochs. These experiments aim to identify the optimal configuration and ensure fair comparisons across models.
@@ 4.3 Results
In Table 6, YOLOv11 [b24] demonstrates the highest overall AP of 0.451, reflecting performance averaged across multiple IoU thresholds ranging from 0.5 to 0.95. Its strong performance on classes such as PET and ECAL, which possess highly distinguishable visual features, suggests its potential for reliable deployment in automated sorting systems. Contrarily, the low AP for Cardboard, likely due to its underrepresentation in the dataset, as shown in Table 2. Augmentation strategies may be critical for improving performance on low-frequency classes.

### v2

@@ 1 Introduction
The exponential growth of the global population and industrial development has increased waste production, making it one of the most pressing environmental challenges of the 21st century. Inefficient waste management systems negatively impact human health, ecosystems, water quality, and the environment at large [24]. As a result, developing innovative and more effective ways to manage and reduce waste is crucial to building a more sustainable future. Solid waste management is a key challenge, with global generation estimated at 2.01 billion tons annually and projected to reach 3.40 billion tons by 2050 under current trends [9].
In most European countries, two principal systems are used for waste sorting. The first is a single-stream system, known as Mechanical-Biological Treatment (MBT), which processes Municipal Solid Waste (MSW). The second is a multi-stream system in which materials are selectively collected through source separation, such as plastics. The main challenge for the single-stream system is that valuable recyclables, particularly plastics, are often incorrectly discarded, limiting their recovery through proper sorting. This loss increases landfill use and consumes raw resources, sabotaging sustainability goals. However, as noted in [4], MBT systems in urban areas can sometimes achieve recovery rates comparable to selective collection. This suggests that with improvements, MBT sorting could be highly impactful. Building on these challenges, deep learning and computer vision offer new opportunities for waste detection and recycling, improving efficiency, reducing reliance on manual labor, and mitigating health risks for workers exposed to hazardous materials. Nevertheless, automatic waste sorting remains challenging because real-world waste streams often contain objects that are deformed, broken, occluded, overlapping, or contaminated [1].
Among the reasons for the lack of automation in waste sorting is the scarcity of publicly available datasets. Existing public datasets typically do not reflect the complex environments found in real-world waste sorting facilities [23, 15, 11, 19, 26, 17, 2, 3, 14], except one contribution [1]. While some companies have developed proprietary datasets for internal use [7, 28, 27], these datasets are not publicly released, restricting the scientific community’s ability to advance research in this area.
@@
The project repository is available at https://github.com/sarainacio/SortWaste.
@@ 2.1 Waste Detection Datasets
Litter Detection. The Garbage in Photos (GINI) dataset [15] contains images of a single class. It was created using the Bing Image Search API and comprises approximately 1,400 images. Similarly, the WADE-AI dataset [11] includes pictures collected from Google Street View, providing a diverse visual set, though it is also limited to a single generic waste classification. Another relevant dataset is TACO (Trash Annotations in Context) [19], introduced by Proença et al., designed for waste detection and segmentation. It contains 1,500 mobile-acquired images annotated into 28 waste categories and covers diverse real-world contexts, including outdoor environments, making it valuable to unconstrained applications. OpenLitterMap [17] is one of the largest public litter datasets, with over 100,000 user-contributed mobile images from around the world, providing diverse scenes and photographic styles that enhance its suitability for real-world litter detection.
@@
Marine Debris. The J-EDI dataset [8] consists of 5,720 images grouped into three classes, created to train deep learning models that enable autonomous underwater vehicles (AUVs) to detect waste in aquatic environments. In addition, AquaTrash, presented in [18] and derived from the TACO dataset, contains 369 multi-class annotated images. This dataset was developed to support AquaVision, a deep learning model aimed at detecting and classifying pollutants in the ocean and along the seashore, contributing to water pollution monitoring and mitigation.
@@
General Waste Objects. The Waste Pictures dataset [26] was constructed using images collected through Google searches. It contains approximately 24,000 images, organized into 34 categories representing a wide variety of waste types. The Taiwan Recycled Waste Dataset (TRWD) [14] comprises 6,233 images featuring multiple objects and is specifically tailored to the unique waste characteristics found in Taiwan. The dataset is divided into six classes.
@@
Recycling Waste. Thung et al. introduced TrashNet [23], a widely used benchmark for waste classification comprising 2,400 images in six categories: metal, paper, plastic, glass, organic, and other, captured against a white background under varying lighting. The WaDaBa dataset [2] focuses exclusively on domestic plastic waste, with 100 plastic objects photographed 40 times each under different conditions, making it particularly relevant for controlled plastic waste identification. Labeled Waste in the Wild (LWW), introduced by Sousa et al. in [3], is a custom dataset containing 1,000 images of waste in food trays captured under real-world conditions. ZeroWaste-f [1] is one of the first high-quality waste datasets collected in real-world conditions. Acquired at a Material Recovery Facility (MRF) in the USA from a high-quality paper sorting line, it is split into three parts for different learning paradigms; the supervised ZeroWaste-f subset contains 4,503 annotated images with four classes. While numerous datasets exist for waste detection tasks, most are not specifically designed for industrial applications, except for the ZeroWaste dataset [1]. Additionally, many datasets consist of images of a single object on a plain white background, which limits the generalizability and effectiveness of models trained on such data when applied to real-world environments. Table 1 summarizes the reviewed datasets, and Figure 2 shows representative examples.
@@
TACO [19] 1500 28 Segmentation Mask R-CNN Litter detection
WADE-AI [11] 1396 1
@@
OpenLitterMap [17] 100000+ 100+ Classification - Litter detection
GINI[15] 1400 1 Classification - Litter detection
@@
AquaTrash [18] 369 4 Detection Faster R-CNN, RetinaNet Marine debris
@@
TRWD [14] 6233 6 Detection YOLOv3 Waste detection
Waste Pictures [26] 23633 34 Classification - Trash objects
@@
TrashNet [23] 2400 6 Classification SVM Recycling waste
WaDaBa [2] 4000 6 Classification - Recycling waste
LWW [3] 1002 19 Detection Faster R-CNN Recycling waste
ZeroWaste-f [1] 4503 4
@@ 2.2 Object Detection Models
Faster R-CNN [22] is a two-stage object detector that extends R-CNN [5] and Fast R-CNN [6] by introducing a Region Proposal Network (RPN), which replaces hand-crafted methods with a fully convolutional network using anchor boxes across scales and aspect ratios. The generated proposals are refined by a Fast R-CNN [6] head for classification and bounding-box regression, and shared convolutional features between the RPN and the detector enable an efficient, end-to-end trainable pipeline that balances accuracy and speed.
Mask R-CNN [6] extends Faster R-CNN [22] by adding a parallel branch for instance segmentation: for each Region of Interest (RoI), it predicts a binary mask alongside the class label and bounding box, enabling joint object detection and pixel-level instance segmentation.
Trident Network [12] addresses the challenge of scale variation in object detection, which often limits the performance of models such as Faster R-CNN [22] and Mask R-CNN [6]. The key idea is to use trident blocks: a multi-branch architecture with shared parameters but different dilation rates. Each branch operates with a distinct receptive field, allowing the network to capture objects at different scales and perform scale-aware detection without relying on external image or feature pyramids. This design improves the performance of two-stage detectors while maintaining the inference complexity.
RetinaNet [13], introduced by Lin et al., is a one-stage detector that reaches accuracy comparable to state-of-the-art two-stage models. Its key contribution is Focal Loss, which mitigates the severe foreground–background imbalance in dense detection. The architecture combines a ResNet–FPN backbone with two subnetworks for anchor classification (using Focal Loss) and bounding box regression. This design substantially closes the accuracy and speed gap between one-stage and two-stage detectors.
YOLO family of detectors [20] formulates object detection as a single-stage regression problem, enabling real-time performance. Successive versions of YOLO, from YOLOv1 to YOLOv10 [21, 25], added features such as multi-scale detection, new activation functions, and anchor-free designs. The YOLOv11 [10] further improves efficiency and accuracy with components like the C3k2 block, SPPF module, and C2PSA mechanism, and extends the framework to instance segmentation, pose estimation, and multi-view recognition.
@@ 3.2 Annotation
The frame rate was downsampled to 5 fps to preserve variability while reducing annotation workload. In total, 18 minutes of video were annotated, yielding 5,396 frames with dirty, deformed, broken, and overlapping objects. Annotations were created over approximately two months using CVAT [16], with each frame requiring, on average, six minutes, depending on scene complexity and object count. This careful annotation process was essential to ensure the quality of the densely labeled dataset. Although there are various types of plastic materials, some are challenging to distinguish using computer vision. Therefore, the annotation focused on simpler and more easily identifiable plastic classes.
@@ 4.1 Implementation Details
We use Faster R-CNN [22], TridentNet [12], and RetinaNet [13], initialized with weights pretrained on the COCO dataset. For each model, only the learning rate and optimizer are tuned, while the batch size is fixed at eight. Training employs early stopping with a patience of 15 epochs or a maximum of 80,000 iterations. For YOLOv11 [10], we adopt a similar tuning strategy, focusing on the learning rate and optimizer, with early stopping defined as 15 epochs of patience or a maximum of 300 epochs. These experiments aim to identify the optimal configuration and ensure fair comparisons across models.
@@ 4.3 Results
In Table 6, YOLOv11 [10] demonstrates the highest overall AP of 0.451, reflecting performance averaged across multiple IoU thresholds ranging from 0.5 to 0.95. Its strong performance on classes such as PET and ECAL, which possess highly distinguishable visual features, suggests its potential for reliable deployment in automated sorting systems. Contrarily, the low AP for Cardboard, likely due to its underrepresentation in the dataset, as shown in Table 2. Augmentation strategies may be critical for improving performance on low-frequency classes.